OpenStudy (anonymous):

2/3 of what number is 16

OpenStudy (anonymous):

put this into an equation. 2/3 times what number? give you 16. let x be that number. Try to solve it now

OpenStudy (confluxepic):

The word "of" translates into a multiplication sign.

OpenStudy (anonymous):

^^ yes that is really important, forgot to mention that.

OpenStudy (jdoe0001):

say the number is "x" so "x" is the whole amount, or 1/1 thus \(\large \begin{array}{ccllll} amount&ration \\\hline\\ 16&\frac{2}{3}\\ x& \frac{1}{1} \end{array}\implies \cfrac{16}{x}=\cfrac{\frac{2}{3}}{\frac{1}{1}}\)